# Maximal variation of linear systems

### v3

@@ Introduction
There is a simple differential criterion for this to happen, namely $H^{0}(T_{X|S})=0$ (§1). However this turns out to be surprisingly difficult to check. We will discuss some cases where maximal variation is expected: hypersurfaces (§2), double coverings of $\mathbb{P}^{n}$ (§3), K3 surfaces and hyperkähler manifolds (§4), and abelian varieties (§5). We prove indeed that the variation is maximal in the latter case, while we have only partial results in the former cases. Finally we discuss some more examples and counter-examples (§6 and 7).
@@ 4. K3 and hyperkähler
For a polarized K3 surface $(X,L)$, it is conjectured that the linear system $\lvert{L}\rvert$ has maximal variation – in fact, it is shown in [D-H, Proposition 5.2] that this would follow from a conjecture of Matsushita on Lagrangian fibrations of hyperkähler manifolds. Here are some known cases:
@@
Let $X$ be a K3 surface, and $L$ an ample line bundle on $X$, with $L^{2}=2g-2$. Then $\lvert{L}\rvert$ has maximal variation in the following cases:
$1)$ $g=2$, or $g=3$ and $(X,L)$ is not hyperelliptic;
$2)$ $g\geq 11$ and the Clifford index of the smooth curves in $\lvert{L}\rvert$ is $\geq 3$ ((this holds in particular when $\operatorname{Pic}(X)=\mathbb{Z}\cdot[L])$;
$3)$ $(X,L)$ is very general;
$4)$ $L=M^{m}$ with $m\geq 2$, for an ample, globally generated line bundle $M$ such that the locus of non-reduced curves in $\lvert{H}\rvert$ has codimension $\geq 3$.
Proof : 1) The case $g=2$ is proved in [D-H, Proposition 5.5], see also Corollary 2. If $g=3$ and $(X,L)$ is not hyperelliptic, $X$ is a smooth quartic surface in $\mathbb{P}^{3}$ and $L=\mathcal{O}_{X}(1)$, so Corollary 1 applies.
2) is proved in [C-D-S, Proposition 2.4]; the last assertion follows from [L].
3) follows from [D-H, Corollary 5.3]; 4) is proved in [D-H, Theorem 1.1].∎
The case of higher-dimensional hyperkähler manifolds turns out to be much easier:
@@ 6. Some more examples
Let $X$ be a surface such that $\Omega^{1}_{X}$ is generated by its global sections. Then any very ample linear system $\lvert{L}\rvert$ has maximal variation.
@@
We observe that $X$ is a minimal surface: if $R$ is a smooth curve in $X$, $\Omega^{1}_{R}$ is a quotient of $\Omega^{1}_{X}$, hence it must be globally generated, which implies $g(R)\geq 1$. Browsing through the classification of surfaces we get the following cases:
$\bullet\ X$ is an abelian surface. This case is settled by Proposition 7.
$\bullet\ X$ admits an elliptic fibration $p:X\rightarrow B$. Since $K_{X}$ is globally generated, it contains a line bundle $p^{*}M$ with $\deg M\geq 1$. On the other hand, a general curve $C\in\lvert{L}\rvert$ cannot be a fiber nor a section of $p$, so $K_{X}\cdot C\geq(p^{*}M)\cdot C>1$.
$\bullet\ X$ is of general type. Then by the index theorem we have $(K_{X}\cdot C)^{2}\geq K_{X}^{2}\cdot C^{2}>1$, hence $K_{X}\cdot C>1$ as required.∎
@@ 7. Some counter-examples
$\bullet$ There are varieties of general type admitting a linear system with no base point and no moduli. Take for instance the Fermat hypersurface $F:\sum\limits_{i=0}^{n}x_{i}^{d}=0$ in $\mathbb{P}^{n}$. For $1\leq r\leq n-2$, consider the rational map $p:F\dasharrow\mathbb{P}^{r}$ given by $p(x_{0},\ldots,x_{n})=\allowbreak(x_{0},x_{1},\ldots,x_{r})$. Let $\hat{F}\rightarrow F$ be the blowing up of the codimension $(r+1)$ subvariety of $F$ given by $x_{0}=\ldots=x_{r}=0$. Then $p$ extends to a morphism $\hat{p}:\hat{F}\rightarrow\mathbb{P}^{r}$ whose general fiber is isomorphic to the Fermat hypersurface of degree $d$ in $\mathbb{P}^{n-r}$. The corresponding linear system $\lvert{\hat{p}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)}\rvert$ has dimension $r$, and no moduli.

### v4

@@ Introduction
There is a simple differential criterion for this to happen, namely $H^{0}(T_{X|S})=0$ (§1). However this turns out to be surprisingly difficult to check. We will discuss some cases where maximal variation is expected: hypersurfaces (§2), double coverings of $\mathbb{P}^{n}$ (§3), K3 surfaces and hyperkähler manifolds (§4), and abelian varieties (§5). We show indeed that the variation is maximal in the last three cases, while we have only partial results in the two first cases. Finally we discuss some more examples and counter-examples (§6 and 7).
@@ 4. K3 and hyperkähler
Any ample linear system on a K3 surface has maximal variation.
This is [Ba, Corollary 2] (which uses [D-H, Proposition 5.2]).∎
The case of higher-dimensional hyperkähler manifolds turns out to be easier:
@@ 6. Some more examples
Let $X$ be a surface of general type such that $\Omega^{1}_{X}$ is generated by its global sections. Then any very ample linear system $\lvert{L}\rvert$ has maximal variation.
@@
We observe that $X$ is a minimal surface: if $R$ is a smooth curve in $X$, $\Omega^{1}_{R}$ is a quotient of $\Omega^{1}_{X}$, hence it must be globally generated, which implies $g(R)\geq 1$. Then by the index theorem we have $(K_{X}\cdot C)^{2}\geq K_{X}^{2}\cdot C^{2}>1$, hence $K_{X}\cdot C>1$ as required.∎
@@ 7. Some counter-examples
$\bullet$ There are varieties of general type admitting a base point free pencil with no moduli. Take for instance the Fermat hypersurface $F:\sum\limits_{i=0}^{n}x_{i}^{d}=0$ in $\mathbb{P}^{n}$. Consider the rational map $p:F\dasharrow\mathbb{P}^{1}$ given by $p(x_{0},\ldots,x_{n})=\allowbreak(x_{0},x_{1})$. Let $\hat{F}\rightarrow F$ be the blowing up of the codimension $2$ subvariety of $F$ given by $x_{0}=x_{1}=0$. Then $p$ extends to a morphism $\hat{p}:\hat{F}\rightarrow\mathbb{P}^{1}$ whose general fiber is isomorphic to the Fermat hypersurface of degree $d$ in $\mathbb{P}^{n-1}$. The pencil $\lvert{\hat{p}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)}\rvert$ has no moduli.